# A Review of Scalable and Privacy-Preserving Multi-Agent Frameworks for Distributed Energy Resource Control

### v1

@@ Abstract
Distributed and decentralized multi-agent systems, distributed energy resources, power and energy systems, privacy preservation
I Introduction
@@
Distributed energy resources (DERs), including solar photovoltaics (PVs), wind turbines, fuel cells, energy storage systems (ESSs), and electric vehicles (EVs), refer to a variety of small-scale energy generation and storage devices that are connected to the electric power grid [1]. They can be controlled individually or in aggregate to provide both grid-level and customer-side benefits, such as providing ancillary services, reducing energy bills, decarbonizing power and energy systems, and enhancing grid resilience [2, 3, 4]. The growth of DERs continues to proliferate, where the global DER management system market is projected to grow from USD 0.42 billion in 2021 to USD 1.33 billion in 2028 at a compound annual growth rate of 18.0% during the 2021-2028 period [5]. The DER market in the U.S. is anticipated to nearly double in capacity from 2022 to 2027, with capital expenditure reaching USD 68 billion per year [6]. The power grid is transitioning towards a DER-populated electricity system where the management of DERs plays an essential role in achieving grid sustainability, resilience, and cybersecurity [7].
Advanced control, optimization, and machine learning theories and tools are essential to fully realize the potential of DERs, especially for achieving scalability in large-scale power systems. These methodologies can assist in solving the DER management problem via generic mathematical formulations with grid objectives and constraints. Broadly, the scalable control of DERs within power systems can be interpreted through a networked multi-agent (we refer to an element of a DER system as an agent) problem where agents can operate in parallel. The unprecedented deployment rates of DERs require scalable management solutions on grid-tied resources to achieve full decarbonization at scale [8]. Besides, building scalable deployment models can accelerate the adoption of the key commercially available but underutilized grid solutions needed to maintain a reliable, safe, and affordable grid [9].
Figure 1: Review structure of scalable and privacy-preserving multi-agent frameworks for DER control.
@@
Another key consideration is privacy protection. Privacy breaches can happen during the processing and transmission of DER data, such as malicious interception of private information during data transmission and the loss of data provenance in the face of dishonest agents [10, 11, 12, 13]. By analyzing load data, adversaries can infer consumers’ lifestyle patterns, personal preferences, and occupancy profiles [14, 15, 16], thereby leading to privacy risks. The threat of privacy leakages targeting the power electric sector, especially DER-populated electric power grids, is escalating in both frequency and complexity. In recent years, a series of stringent privacy protection laws have come into effect to increase protections for consumers’ personal data. These include the strongest privacy and security law [17], European Union’s General Data Protection Regulation, effective in 2018, the U.S.’s first privacy law California Consumer Privacy Act [18], also effective in 2018, the Virginia’s Consumer Data Protection Act [19], effective in 2023, and the most recent Texas Data Privacy and Security Act [20], effective in 2024, and other similar laws [21]. The increased privacy awareness in legislation is driving privacy protection measures and standards in a broad spectrum of cyber and physical systems, such as health care systems [22], wireless networks [23], and power systems in this regard [24]. Therefore, achieving scalability and protecting privacy are two key factors in deploying advanced multi-agent frameworks to optimize DER operations.
@@
A number of existing reviews have underlined the importance of DER control, along with arising related privacy and cybersecurity concerns [25, 26, 27, 28, 29, 30, 31]. Zografopoulos et al. [25] point out cybersecurity issues in DER control problems caused by adversarial capabilities and objectives, and review both DER protocol-level and DER device-level vulnerabilities, attacks, impacts, and mitigations. Ghiasi et al. [26] provide a comprehensive review of cyberattacks and defense mechanisms for smart grid energy systems. In [27], the authors review privacy-preserving schemes for smart grid applications, addressing privacy leakages from key-based, data-based, impersonation-based, and physical-based attacks. Considering the increasing penetration of renewable energy, Tuyen et al. [28] survey state-of-the-art detection and mitigation techniques, with a focus on the system structure and vulnerabilities of typical inverter-based power systems integrated with DERs. In [29], the authors review standards, protocols, and constraints and provide recommendations for mitigating cyber-attacks in cyber-physical power systems. Liu et al. [30] survey developments on enhancing cyber-resiliency of DER-based smart grids, including threat modeling, risk assessment, and defense-in-depth strategies. In [31], Cardenas et al. underscore the need for privacy-aware solutions and discuss grid-related digital privacy risks. Notably, the aforementioned reviews provide different examinations of controlling DERs from multi-faceted perspectives. However, an interdisciplinary review of orchestrated scalable and privacy-preserving solutions is important to deploy advanced multi-agent DER control strategies in practice.
@@
Motivated by the proliferation of recent research outcomes on DER controls, this paper reviews state-of-the-art techniques for designing scalable and privacy-preserving multi-agent frameworks and their applications to DER control problems. Fig. 1 provides an overview of the review structure. We first survey scalable multi-agent frameworks based on contemporary distributed and decentralized information exchange structures, and then review integrated privacy preservation techniques from the perspective of privacy-aware multi-agent computing frameworks.
@@
    We give a systematic review of deploying multi-agent frameworks for DER control in power systems regarding multi-agent-based problem formulation, scalable solutions, and privacy preservation techniques.
@@
    We survey state-of-art scalable algorithms within multi-agent frameworks based on distributed and decentralized information exchange structures and review representative works for DER control problems. Moreover, we identify internal, external, and hierarchical types of adversaries in multi-agent systems that can compromise the system’s privacy and security.
@@
Figure 2: Three typical information exchange structures of networked multi-agent systems for managing DERs in power systems: (a) Centralized information exchange that relies on a system operator to collect information from all agents, process it, and then send control commands to each agent; (b) Distributed structure that allows agents to operate independently, interact with coordinator/environment, and communicate with each other over a network; and (c) Decentralized structures that is similar to a Distributed structure, but without peer-to-peer communications.
In the rest of this paper, Section II gives an overview of multi-agent systems and their applications for DER control in power systems. Section III details the large-scale DER control problem by constructing a multi-agent optimization model. Section IV surveys predominantly scalable methods for addressing the multi-agent problem and summarizes privacy issues associated with these scalable approaches. Section V reviews privacy preservation techniques for scalable multi-agent-based DER control. Section VI extrapolates new approaches for future scalable, privacy-aware, and cybersecure pathways to unlock the full potential of DERs. Section VII concludes the paper.
II Overview of Multi-Agent Systems
The cooperative grid-level objectives support the achievement of system-wide goals, such as achieving overnight valley filling [33, 34, 35], minimizing power lines losses [36, 37, 38], reducing the emission of pollutants [39, 40], etc. The competitive DER-level objectives aim at maximizing the benefits of DERs, such as bidding in the electricity market [41], reducing energy costs for consumers [42, 43], and minimizing battery degradation costs [44]. The networked constraints can include the nodal voltage deviations and power flow constraints [45], which are coupled through the power network topology. The individual DER constraints can include battery charging/discharging rate [46], the capacity of generators [47], solar power availability for PV curtailment [48], etc.
The information exchange model defines the computing and communication structure for solving networked multi-agent problems. Various models have been developed to facilitate control, optimization, and learning in these multi-agent systems [35, 49, 50, 51, 52, 53, 54]. In summary, these models can be classified into centralized, distributed, and decentralized structures, as shown in Fig. 2. Following this classification, this paper reviews scalable multi-agent frameworks within distributed and decentralized structures, addressing the interests of different stakeholders when operating DER-populated power systems. The system operator (SO) (e.g., distribution or transmission SO) functions as a central authority and can provide instructions on coordinating and controlling the power system operations. To clarify, we refer to the coordinator as a central entity that is needed solely for the coordination of signals rather than for directly controlling any agent.
In a centralized setting, the SO manages the entire power system operation by collecting agent and network information, processing it, and then sending control commands to all agents [35]. Therefore, the DER control problem is solved in a central control room where the SO makes strategic decisions on achieving grid-level and/or DER-level objectives, while agents simply follow the SO’s commands. Centralized approaches are easy to implement and can often obtain globally optimized solutions. However, they suffer from drawbacks caused by 1) computing and communication overheads imposed on the SO, 2) compromised data privacy and security, and 3) vulnerability during cyber and physical contingencies. Due to the heavy dependence on the SO, centralized information exchange is more effective for small-scale cases with a modest agent population size.
Similarly, decentralized approaches also achieve scalability by allocating central computing loads to each local agent, but with an emphasis on eliminating peer-to-peer communications. In decentralized information exchange structures, agents make decisions independently in a possible networked environment without communicating with each other. However, they may interact with the environment directly or rely on the assistance of a coordinator. In a typical decentralized framework, such as primal-dual-based algorithms [52, 53], agents and the coordinator iteratively update the primal variable (decision variable) and the dual variable (Lagrange multiplier), respectively. Owing to the outstanding scalability, distributed and decentralized multi-agent frameworks are well suited for large-scale DER control problems.
@@ II Overview of Multi-Agent Systems
III Preliminaries on Multi-Agent-Based DER Control
The DER control problem can be described as a multi-agent optimization model that defines the objectives, constraints, and decision variables to optimize the power grid operations. The decision-making can be achieved by solving the formulated multi-agent problem with control, optimization, and learning-based methods. In this section, we present a general multi-agent problem formulation and then delve into the detailed objectives and constraints for multi-agent-based DER control problems.
III-A DERs in Power Systems
With the rising reliance on DERs, the energy transition is changing how loads and generations need to be managed for homes and utilities. DERs can provide cheaper, cleaner, and more accessible electricity supplies by connecting at various points within a power system, such as distribution and transmission networks, substations, and behind the meter. Evidently, the fast adoption of DERs provides multifarious grid services, such as voltage regulation, demand response, and enhanced customer satisfaction [68]. Furthermore, past grid failures demonstrate the vulnerability of power grids under extreme climate events [69, 70, 71], operational failures [72, 73], and cyber and physical attacks [74, 75], therefore compelling SOs, prosumers, and consumers to rely more on DERs to enhance grid resilience, both individually and in aggregate.
To unlock the full potential of DERs, this paper focuses on two key technical challenges in optimizing DER-populated power systems, i.e., scalability and privacy. The old model of centralized electrical supply is no longer the sole reality, massive DERs with varying attributes require a scalable management plan. Furthermore, to capture the DER market and enhance customer engagement, privacy-preserving decision-support tools need to be developed, integrated, and tested. These tools, in turn, can optimize customers’ electrical energy services. The synergy of scalability and privacy protection is becoming a trending research topic among the control, optimization, learning, and power communities. Fig. 3 shows the conceptualization of a privacy-preserving DER management system, where the prosumers, customers, aggregators, utilities, and distribution network/system operators (DNO/DSO) collaborate to achieve grid-level objectives and customer-side goals.
@@ III-A DERs in Power Systems
Problem (1) can be broadly applied to a variety of power system applications, with goals such as grid modernization, decarbonization, and resilience. In the following section, we provide details on multi-agent-based DER control using the formulation of Problem (1).
III-B Network Models, Objectives, and Constraints
III-B1 Network Models
@@
The power flow model is built upon power network topology, loads, and generations. DERs can function as flexible loads or generation units in power distribution and transmission networks. We next present the control of DERs in distribution systems using the nonlinear DistFlow branch model [76]. Consider a radial distribution network described by a connected graph $G=\{\mathcal{N},\mathcal{E}\}$, where $\mathcal{N}=\{0,1,\ldots,n\}$ denotes a set of nodes/buses, $\mathcal{E}\subset\mathcal{N}\times\mathcal{N}$ denotes a set of directed edges/lines. The network is tree-structured where Node 0 serves as the slack bus and maintains a constant voltage magnitude of $V_{0}$.

### v2

# A Review of Scalable and Privacy-Preserving Multi-Agent Frameworks for Distributed Energy Resources

@@ Abstract
Decentralized multi-agent systems , distributed energy resources , power systems , privacy preservation , cyber-physical system security
journal: Advances in Applied Energy
[…]
[label1]organization=Department of Electrical and Computer Engineering, Texas A&\&&M University,city=College Station, postcode=77843, state=TX, country=USA
[…]
[label2]organization=Department of Electrical and Computer Engineering, Princeton University,city=Princeton, postcode=08544, state=NJ, country=USA
[…]
[label3]organization=Department of Electrical and Computer Engineering, University of Utah,city=Salt Lake City, postcode=84112, state=UT, country=USA
1 Introduction
1.1 Significance of DERs
@@
Distributed energy resources (DERs), including solar photovoltaics (PVs), wind turbines, fuel cells, energy storage systems (ESSs), and electric vehicles (EVs), refer to a variety of small-scale energy generation and storage devices that are connected to the electric power grid [1]. DERs offer substantial flexibility to power systems at both the grid and customer levels, such as providing ancillary services [2], lowering energy costs [3], decarbonizing power systems [4, 5], and enhancing grid resilience [6]. Because of these benefits, the power grid is transitioning toward an increasingly DER-rich electricity system, where the management of DERs is crucial for supporting the integration of renewable energy, enhancing grid resilience, and improving overall energy efficiency [7]. As the energy transition accelerates, the growing dependence on DERs is reshaping how loads and generation sources are managed for homes and utilities, offering cheaper, cleaner, and more accessible electricity supplies. Additionally, past grid failures from extreme weather events, operational breakdowns, and cyber-physical attacks [8, 9, 10, 11, 12, 13, 14] have underscored the need for system operators (SOs), prosumers, and consumers to increasingly rely on DERs, both individually and in aggregate, to bolster grid resilience. The global DER management system market is expected to expand significantly, projected to increase from USD 0.42 billion in 2021 to USD 1.33 billion by 2028, reflecting a compound annual growth rate of 18.0% during this period [15]. In the U.S., the DER market is anticipated to nearly double in capacity from 2022 to 2027, with capital expenditures reaching USD 68 billion per year [16]. Hence, the importance of DERs in power systems can be seen by their rapid growth and their pivotal role in modernizing the grid.
1.2 Major DER Control Challenges
1.2.1 Scalability Issues
Traditionally, the power grid is managed in a centralized manner, with a single control problem formulated using data collected from the wide-area power grid network. However, the rapid increase of DERs at various scales makes it challenging to solve such large-scale centralized problems, due to the growing complexity, amount of heterogeneous data, and high computing costs. The high deployment of DERs requires scalable management solutions for grid-tied resources to meet grid objectives, such as full decarbonization at scale [4]. Besides, building scalable deployment models can accelerate the adoption of commercially available grid solutions [17]. Broadly, the scalable control of DERs within power systems can be interpreted through a networked multi-agent (we refer to an element of a DER system as an agent) problem where agents can operate in parallel. To fully realize the potential of DERs to help achieve greater sustainability and resilience in power systems, it is essential to develop scalable multi-agent frameworks incorporating advanced control, optimization, and machine learning theories and tools. These advanced techniques can help solve the DER management problems efficiently via generic mathematical formulations in a scalable fashion that incorporates grid objectives and constraints. For example, in [18], Radhakrishnan and Srinivasan propose a multi-agent-based DER management system to perform optimal management of DERs. Wang et al. [19] propose a data-driven multi-agent power grid control scheme using a deep reinforcement learning (RL) method for autonomous voltage control. In [20], Fan et al. develop a multi-agent distributed control framework to realize the distributed optimal generation control with multiple DERs in real-time. All these developments highlight the benefits of developing multi-agent frameworks with advanced techniques for DER management to ensure the optimality, safety, and security of power grid operations.
1.2.2 Privacy Threats
@@
Another key consideration is privacy preservation/protection. Privacy breaches can happen during the processing and transmission of DER data, such as malicious interception of private information during data transmission and the loss of data provenance in the face of dishonest agents [21, 22, 23, 24]. By analyzing load data, adversaries can infer consumers’ lifestyle patterns, personal preferences, and occupancy profiles [25, 26, 27], thereby leading to privacy risks. The threat of privacy leakages targeting the power electric sector, especially DER-rich electric power grids, is escalating in both frequency and complexity. In recent years, a series of stringent privacy protection laws have come into effect to increase protections for consumers’ personal data. These include the strongest privacy and security law [28], European Union’s General Data Protection Regulation, effective in 2018, the U.S.’s first privacy law California Consumer Privacy Act [29], also effective in 2018, the Virginia’s Consumer Data Protection Act [30], effective in 2023, and the most recent Texas Data Privacy and Security Act [31], effective in 2024, and other similar laws [32]. The increased privacy awareness in legislation is driving privacy protection measures and standards in a broad spectrum of cyber and physical systems, such as health care systems [33], wireless networks [34], and power systems in this regard [35]. For the cybersecurity of smart grids, National Institute of Standards & Technology (NIST) has established the Guidelines for Smart Grid Cybersecurity to develop effective cybersecurity strategies for protecting the privacy of smart grid-related data and for securing the communication networks [36]. The U.S. Department of Energy, in collaboration with industry stakeholders, launched the DataGuard Energy Data Privacy Program to develop privacy concepts and principles for grid customers [37]. Europe Commission has introduced the Data Protection Impact Assessment template for smart grid and smart metering systems, aimed at assessing risks and developing countermeasures to ensure the protection of personal data [38]. Additionally, the International Electrotechnical Commission (IEC) has specified cybersecurity requirements for smart grids through a series of standards in IEC 62351, addressing secure data transfer, prevention of eavesdropping, and intrusion detection [39]. Therefore, eliminating privacy and security concerns is critical for deploying advanced multi-agent frameworks to optimize DER operations.
1.3 Motivation and Contributions
@@
To unlock the full potential of DERs, this paper focuses on two key technical challenges in optimizing DER-rich power systems, i.e., scalability and privacy. The old model of centralized electrical supply is no longer the sole reality, massive DERs with varying attributes require a scalable management plan. Furthermore, to capture the DER market and enhance customer engagement, privacy-preserving decision-support tools need to be developed, integrated, and tested. These tools, in turn, can optimize customers’ electrical energy services. The synergy of scalability and privacy protection is becoming a trending research topic among the control, optimization, learning, and power communities. A number of existing reviews have underlined the importance of DER control, along with arising related privacy and cybersecurity concerns [40, 41, 42, 43, 44, 45, 46]. Zografopoulos et al. [40] point out cybersecurity issues in DER control problems caused by adversarial capabilities and objectives, and review both DER protocol-level and DER device-level vulnerabilities, attacks, impacts, and mitigations. Ghiasi et al. [41] provide a comprehensive review of cyberattacks and defense mechanisms for smart grid energy systems. In [42], the authors review privacy-preserving schemes for smart grid applications, addressing privacy leakages from key-based, data-based, impersonation-based, and physical-based attacks. Considering the increasing penetration of renewable energy, Tuyen et al. [43] survey state-of-the-art detection and mitigation techniques, with a focus on the system structure and vulnerabilities of typical inverter-based power systems integrated with DERs. In [44], the authors review standards, protocols, and constraints and provide recommendations for mitigating cyber-attacks in cyber-physical power systems. Liu et al. [45] survey developments on enhancing cyber-resiliency of DER-based smart grids, including threat modeling, risk assessment, and defense-in-depth strategies. In [46], Cardenas et al. underscore the need for privacy-aware solutions and discuss grid-related digital privacy risks. Notably, the aforementioned reviews provide different examinations of controlling DERs from multi-faceted perspectives. However, an interdisciplinary review of orchestrated scalable and privacy-preserving solutions is important to deploy advanced multi-agent DER control strategies in practice.
@@
Motivated by the proliferation of recent research outcomes on DER controls, this paper reviews state-of-the-art techniques for designing scalable and privacy-preserving multi-agent frameworks and their applications to DER control problems.
Figure 1: Review structure of scalable and privacy-preserving multi-agent frameworks for DERs.
Fig. 1 provides an overview of the review structure. We first survey scalable multi-agent frameworks based on contemporary distributed and decentralized information exchange structures, and then review integrated privacy preservation techniques from the perspective of privacy-aware multi-agent computing frameworks.
@@
    We conduct and present a systematic review of deploying multi-agent frameworks for DER control in power systems regarding multi-agent-based problem formulation, scalable solutions, and privacy preservation techniques.
@@
    We survey state-of-art scalable algorithms within multi-agent frameworks based on distributed and decentralized information exchange structures, and we review representative works for DER control problems. Moreover, we identify internal, external, and hierarchical types of adversaries/threats in multi-agent systems that can compromise the system’s privacy and security.
@@
In the rest of this paper, Section 2 provides an overview of multi-agent systems for DER control applications in power systems, detailing the corresponding multi-agent optimization model and the information exchange structures. Section 3 surveys predominantly scalable methods for addressing the multi-agent problem and summarizes privacy issues associated with these scalable approaches. Section 4 reviews privacy preservation techniques for scalable multi-agent-based DER control. Section 5 extrapolates new approaches for future scalable, privacy-aware, and cybersecure pathways to unlock the full potential of DERs. Section 6 concludes the paper.
2 Multi-Agent Systems of DERs
The DER control problem can be described using a multi-agent optimization model that defines the objectives, constraints, and decision variables to optimize the power grid operations. The optimal decision-making can be achieved by solving the formulated multi-agent problem with control, optimization, and learning-based methods. In this section, we present a general multi-agent problem formulation and then delve into the detailed objectives and constraints for multi-agent-based DER control problems.
2.1 Multi-Agent-Based Control of DERs
The cooperative grid-level objectives support the achievement of system-wide goals, such as achieving overnight valley filling [48, 49, 50], minimizing power lines losses [51, 52, 53], reducing the emission of pollutants [54, 55], etc. The competitive DER-level objectives aim at maximizing the benefits of DERs, such as bidding in the electricity market [56], reducing energy costs for consumers [57, 58], and minimizing battery degradation costs [59]. The networked constraints can include the nodal voltage deviations and power flow constraints [60, 61], which are coupled through the power network topology. The individual DER constraints can include battery charging/discharging rate [62], the capacity of generators [63], solar power availability for PV curtailment [64], etc. Besides, individual DER constraints often need to account for thermal dynamics, which are coupled with the operation of heating, ventilation, and air conditioning (HVAC) systems [65, 66].
The information exchange model defines the computing and communication structure for solving networked multi-agent problems. Various models have been developed to facilitate control, optimization, and learning in these multi-agent systems [50, 67, 68, 69, 70, 71, 72]. In summary, these models can be classified into centralized, distributed, and decentralized structures, as shown in Fig. 2. Multi-agent frameworks also include hierarchical information exchange structures that combine centralized and distributed computing schemes, leveraging the strengths of both approaches to achieve both global oversight and local autonomy. In a hierarchical framework, a centralized layer typically oversees high-level decision-making and coordination by utilizing global optimization to guide agents toward optimal solutions. On the other hand, the distributed layer allows individual agents to operate autonomously, handling localized tasks and reacting to dynamic environmental changes in real-time. Following this classification, this paper reviews scalable multi-agent frameworks within distributed and decentralized structures, addressing the interests of different stakeholders when operating DER-rich power systems. The system operator (SO) (e.g., distribution or transmission SO) functions as a central authority and can provide instructions on coordinating and controlling the power system operations. To clarify, we refer to the coordinator as a central entity that is needed solely for the coordination of signals rather than for directly controlling any agent.
In a centralized setting, the SO manages the entire power system operation by collecting agent and network information, processing it, and sending control commands to all agents [50]. Therefore, the DER control problem is solved centrally where the SO makes strategic decisions on achieving grid-level and/or DER-level objectives, while agents simply follow the SO’s commands. Centralized approaches are easy to implement and can often obtain globally optimized solutions. However, they suffer from drawbacks caused by 1) computing and communication overheads imposed on the SO, 2) compromised data privacy and security, and 3) vulnerability during cyber and physical contingencies. Due to the heavy dependence on the SO, centralized information exchange is more effective for small-scale cases with a modest agent population size.
Similarly, decentralized approaches also achieve scalability by allocating central computing loads to each local agent, but with an emphasis on eliminating agent-to-agent communications. In decentralized information exchange structures, agents make decisions independently in a possible networked environment without communicating with each other. Compared to distributed structures, decentralized approaches eliminate agent-to-agent communication, thereby reducing privacy risks associated with direct information exchange between agents. However, agents in decentralized structures are often required to interact directly either with the environment or rely on the assistance of a coordinator, both of which can lead to private information leakage, such as in the presence of honest-but-curious agents, external eavesdroppers, and/or the coordinator. In a typical decentralized framework, such as primal-dual-based algorithms [70, 71], agents and the coordinator iteratively update the primal variable (decision variable) and the dual variable (Lagrange multiplier), respectively. Owing to the outstanding scalability, distributed and decentralized multi-agent frameworks are well suited for large-scale DER control problems.
@@ 2.1 Multi-Agent-Based Control of DERs
Figure 2: Three typical information exchange structures of networked multi-agent systems for managing DERs in power systems: (a) Centralized information exchange that relies on a system operator to collect information from all agents, process it, and then send control commands to each agent; (b) Distributed structure that allows agents to operate independently, interact with coordinator/environment, and communicate with each other over a network; and (c) Decentralized structures that is similar to a Distributed structure, but without agent-to-agent communications.
@@
2.2 General Problem Formulation
@@
Problem (1) can be broadly applied to a variety of power system applications, with goals such as grid modernization, decarbonization, and resilience.
Fig. 3 shows the conceptualization of a privacy-preserving DER management system, where the prosumers, customers, aggregators, utilities, and distribution network/system operators (DNO/DSO) collaborate to achieve grid-level objectives and customer-side goals. In the following section, we provide details on multi-agent-based DER control using the formulation of Problem (1).
2.3 Network Models, Objectives, and Constraints
2.3.1 Network Models
@@
The power flow model is built upon power network topology, loads, and generation sources. DERs can function as flexible loads or generation units in power distribution and transmission networks. We next present the control of DERs in distribution systems using the nonlinear DistFlow branch model [86]. Consider a radial distribution network described by a connected graph $G=\{\mathcal{N},\mathcal{E}\}$, where $\mathcal{N}=\{0,1,\ldots,n\}$ denotes a set of nodes/buses, $\mathcal{E}\subset\mathcal{N}\times\mathcal{N}$ denotes a set of directed edges/lines. The network is tree-structured where Node 0 serves as the slack bus and $V_{0}$ denotes the constant voltage magnitude of Node 0.